OpenStudy (anonymous):

The total amount of weight on an airplane cannot exceed 2,500 pounds. The plane is carrying 1,750 pounds of luggage. The average weight of each passenger on the plane is 150 pounds. Set up and solve an inequality to find the number of passengers, p, the plane could carry.

OpenStudy (anonymous):

1750 + 150p = 2500

OpenStudy (anonymous):

Then 750 / 150 which equals so 5 passengers.

OpenStudy (anonymous):

thnx!

OpenStudy (anonymous):

Can I have my medal????

OpenStudy (anonymous):

lol @jasonArnold

OpenStudy (anonymous):

For the people that had answer chooses to this question: A- 150p+1750=<2500; p=<5 B- 150+1750p=<2500; p=<1.3 C- 150p-1750+<2500; p=< 28.3 D- 150p+2500=>1750; p=>-5 So the Answer that he said is A